# Learning Decision-Focused Uncertainty Sets in Robust Optimization

### v4

@@ 4 Probabilistic guarantees of constraint satisfaction
In the previous section, we have proved convergence to a feasible solution of the sampled problem (10), which does not exactly correspond to the probabilistic guarantees of constraint satisfaction (7). To arrive at our desired guarantees, where the inner probability is with respect to the unknown joint distribution ${\mathbf{P}}_{(u,y)}$, we need to analyze the generalization behavior of the robust solutions on out-of-sample data. We begin with a boundedness assumption on the function inside the expectation from the $\mathop{\bf CVaR}$ definition, which appears in (8).
@@
The function $g(x(\theta,y),u,y)-\alpha)_{+}/\eta+\alpha$ maps to values in $(-c^{\rm min},c^{\rm max})$, where $0\leq c^{\rm min},c^{\rm max}<\infty$, for all $\theta,\alpha,u,y$ in its support.
Let $C=c^{\rm min}+c^{\rm max}$ and define the normalized function
$\psi_{z}(u,y)=\frac{1}{C}\left(c^{\rm min}+\frac{(g(x,u,y)-\alpha)_{+}}{\eta}+\alpha\right),$ (13)
where $z=(x,\alpha)\in\mathcal{Z}\subset{\mbox{\bf R}}^{n+1}$. Note that when Assumption 3.4 holds, we have $\mathcal{Z}=\mathcal{X}\times\mathcal{D}_{\alpha}$ being a compact set. The function $\psi_{z}$ then maps to values in $(0,1)$, and we consider the entire function class
$\Psi=\{\psi_{z}\}_{z\in\mathcal{Z}},$
@@
$\sup_{Q}N(\delta,\Psi,L_{2}(Q))\leq\left(\frac{K}{\delta}\right)^{V},$ (14)
for every $0<\delta<K$. Then, when Algorithm 1 converges to an $\epsilon$-stationary point of (10), we have the finite sample probabilistic guarantee
${\mathbf{P}}^{N\times J}\left({\mathbf{P}}_{(u,y)}(g(x,u,y)\leq 0)\geq 1-\eta\quad\forall x\in\mathcal{X}\right)\geq 1-\beta,$ (15)
where $\beta=(D\sqrt{N}C\tau/\sqrt{V})^{V}\exp(-2NC^{2}\tau^{2})$, $D$ is a constant that depends only on $K$, and $\tau\leq(c^{\min}+\kappa+\epsilon)/C$.
@@
${\mathbf{P}}^{N\times J}\left(\sup_{\psi_{z}\in\Psi}\left\|\frac{1}{NJ}\sum_{j=1}^{J}\sum_{i=1}^{N}\psi_{z}(u^{i},y^{j})-{\mbox{\bf E}}_{(u,y)}[\psi_{z}(u,y)]\right\|\leq\tau\right)\geq 1-\left(\frac{D\sqrt{N}C\tau}{\sqrt{V}}\right)^{V}\exp(-2NC^{2}\tau^{2}).$
By the convergence of Algorithm 1 to an $\epsilon$-stationary point of (10), we have that the empirical $\mathop{\bf CVaR}$ will be close to $\kappa$, which, given equation (13), means
$\frac{1}{NJ}\sum_{j=1}^{J}\sum_{i=1}^{N}\psi_{z}(u^{i},y^{j})\leq(1/C)(c^{\min}+\kappa+\epsilon).$
Therefore, we have
${\mathbf{P}}^{N\times J}\left(\sup_{\psi_{z}\in\Psi}{\mbox{\bf E}}_{(u,y)}[\psi_{z}(u,y)]\leq(1/C)(c^{\min}+\kappa+\epsilon)+\tau\right)\geq 1-\beta,$
@@
${\mathbf{P}}^{N\times J}\left(\mathop{\bf CVaR}(g(x,u,y),\eta)\leq 0\quad\forall x\in\mathcal{X}\right)\geq 1-\beta,$
which implies
${\mathbf{P}}^{N\times J}\left({\mathbf{P}}_{(u,y)}(g(x,u,y)\leq 0)\geq 1-\eta\quad\forall x\in\mathcal{X}\right)\geq 1-\beta.$ (16)
As this probabilistic guarantee holds for all $x$ within the bounded set $\mathcal{X}$, it also holds for our data-driven solutions $x(\theta,y)$.
Corollary 4.1.1.
Suppose Assumption 3.4 holds. For all $\theta^{0}\in\mathcal{D}_{\theta}$, $s^{0}\in\mathcal{D}_{s}$, and $y\in\mathcal{D}_{y}$, Theorem 4.1 implies the probabilistic guarantee (7) for the data-driven solution $x(\theta^{\star},y)$, i.e., ${\mathbf{P}}^{N\times J}\left\{{\mathbf{P}}_{(u,y)}(g(x(\theta^{\star},y),u,y)\leq 0)\geq 1-\eta\right\}\geq 1-\beta.$
We note that the covering number requirement holds when $\Psi$ is a class of Lipschitz functions, which is implied when the constraint function $g$ is Lipschitz. For instance, when $g$ is affine or a maximum of affine functions in variable $x$, it is Lipschitz in $x$ for all $u\in\mathcal{D}_{u}$ and $y\in\mathcal{D}_{y}$,  i.e.,
$|g(x,u,y)-g(x^{\prime},u,y)|\leq\bar{L}\|x-x^{\prime}\|_{2}\quad\forall x,x^{\prime}\in\mathcal{X},~{}u\in\mathcal{D}_{u},~{}y\in\mathcal{D}_{y},$
with some Lipschitz constant $\bar{L}$. Then, $\psi_{z}\in\Psi$ is Lipschitz in $z=(x,\alpha)$ with Lipschitz constant $L\leq(1/C)((\bar{L}+1)/\eta+1)$. We can now derive the following covering number condition.
@@
When $\psi_{z}$ is $L$-Lipschitz in $z$ for all $u$ and $y$, we can apply standard bounds on the covering number of Lipschitz losses (van der Vaart and Wellner,, 1996; Pollard,, 1890; Namkoong and Duchi,, 2017). In this case, the covering number condition (14) reduces to
$\sup_{Q}N(\delta,\Psi,L_{2}(Q))\leq N(\delta/L,\mathcal{Z},\|\cdot\|_{2})\leq\left(1+\frac{(x^{\rm max}+\alpha^{\rm max})L}{\delta}\right)^{n+1}\leq\left(\frac{K}{\delta}\right)^{n+1},$
for every $0<\delta<K$, where $K=\delta+(x^{\rm max}+\alpha^{\rm max})L$. Then, $\beta=(D\sqrt{N}C\tau/\sqrt{n+1})^{n+1}\exp(-2NC^{2}\tau^{2})$.
@@ 5 Examples
We run the experiments on the Princeton Institute for Computational Science and Engineering (PICSciE) facility with 20 parallel 2.4 GHz Skylake cores. We solve the inner optimization problems (1) with the Clarabel optimizer with default settings. The learning procedure relies on the Python package cvxpylayers (Agrawal et al., 2019a, ), and the code to replicate the experiments can be accessed at
@@
For each experiment, we compare our method with a RO and a DRO benchmark, given below.
Learning for robust optimization (LRO) (our method).
We apply the training method given in Section 3.2, for a variety of $\eta$ values, to achieve various empirical probabilities of constraint violation $\hat{\eta}$. The uncertainty set can be RO or DRO based. Default training parameters are given in Table 2.
LRO + Fine Tuning (our method).
Upon completion of the LRO training procedure, we also fine tune the size of the uncertainty set using grid search to precisely achieve a target in-sample empirical probability of constraint violation.
Mean-Variance RO.
We perform mean-variance RO, where the shape of the uncertainty set is determined by the empirical mean and variance of the data (Ben-Tal and Nemirovski,, 2000). We use grid search over a the set size to achieve various empirical probabilities of constraint violation $\hat{\eta}$.
Wasserstein DRO.
We perform Wasserstein-1 DRO (Esfahani and Kuhn,, 2018), where the uncertainty set is constructed around individual data points. We again use grid search over the set size to achieve various empirical probabilities of constraint violation $\hat{\eta}$.
For each setting, we repeat the experiment 20 times, each with a new independent dataset. For each method, we evaluate the performance metrics in Table 1.
Table 1: Performance metrics in numerical examples, averaged over 20 runs.
Name Description Obj. Out-of-sample objective value estimate for problem (8) $\hat{\eta}$ Out-of-sample empirical probability of constraint violation $\hat{\beta}$ Out-of-sample estimated confidence, given as ${\mathbf{P}}(\hat{\eta}>\eta)$, where the probability is over the testing data ${\rm time}$ Average solution time of the final robust problem after training
Table 2: Default hyperparameter values for Algorithm 1. We set the initial uncertainty set parameter values $(A^{0},b^{0})=(\hat{\Sigma}^{1/2},\hat{\mu})$, depending on the empirical covariance and mean of the training data.
$\theta^{0}$ $\alpha^{0}$ $s^{0}$ $\lambda^{0}$ $\mu^{0}$ $\kappa$ $\sigma$ $\tau$ $k_{\text{max}}$ $t_{\text{max}}$ $\eta$ $|B|$
$(A^{0},b^{0})$ 0 0 1 1 $-0.01$ 1.015 0.95 1000 100 $[0,0.3]$ 30
@@ 5.1 Portfolio optimization
We consider the portfolio management problem where we select a portfolio $x\in{\mbox{\bf R}}^{n}$ of stocks to maximize $u^{T}x$, where $u\in{\mbox{\bf R}}^{n}$ are the uncertain asset returns. The trading costs are proportional to the deviations from our nominal holdings $x^{\text{nom}}\in{\mbox{\bf R}}^{n}$, i.e., $\lambda^{\rm trade}\|x-x^{\text{nom}}\|_{1}$ with $\lambda^{\rm trade}>0$. We seek to minimize the objective $-u^{T}x+\lambda^{\rm trade}\|x-x^{\text{nom}}\|_{1}.$ We parametralize the problem with the family parameter $y=x^{\text{nom}}$, and solve the RO problem
$\begin{array}[]{ll}\mbox{minimize}&t+\lambda^{\text{trade}}\|x-x^{\text{nom}}\|_{1}\\ \mbox{subject to}&-u^{T}x\leq t\quad\forall u\in{\mathcal{U}}(\theta)\\ &\mathbf{1}^{T}x=1,\quad x\geq 0,\end{array}$
where the uncertain parameter $u$ stands for the unknown returns.
@@
We consider the case where $n=5$ and $n=10$ stocks. We let $\lambda^{\text{trade}}=0.2$. The uncertain returns $u$ is sampled from a multivariate normal distribution $N(\mu,\Sigma)$, where $\mu$ is generated uniformly on $[0.5,1]^{n}$, and $\Sigma$ is generated as $0.1FF^{T}$, where $F$ is generated on $[0,1]^{n\times 2}$. The nominal holdings $\bar{x}^{\text{nom}}$ is generated from a Dirichlet distribution with parameter values equal to $\mu$. We consider 10 instances of $x^{\text{nom}}$, distributed normally around $\bar{x}^{\text{nom}}$, with a variance of $0.05$. Each sample is then projected to be non-negative, and normalized to sum to 1. The total training sample size is $N=500$, with a testing sample size of $20000$.
Uncertainty set.
We learn both a RO and a DRO uncertainty set. For RO, the set is ellipsoidal with with norm $p=2$, and for DRO, we train the shape of a mean robust optimization (MRO) set with $K=5$ and power $p=1$. This set is closely related to the Wasserstein type-1 DRO set. See Appendix A for details on the MRO uncertainty set. By choosing $K\ll N$, we speed up the training procedure with minimal performance degradation compared to training with Wasserstein DRO (Wang et al.,, 2022). For both $n=5$ and $n=10$, we learn uncertainty sets for $\eta\in\eta^{\text{list}}=(0.01,0.03,0.05,0.08,0.10,0.15,0.2,0.3)$.
@@ Results.
In Table 3, we compare the performance of our LRO methods with mean-variance RO and Wasserstein DRO, and note the desirable trade-off given by our learned sets.
[…]
In Figure 3, for the 20 runs of the experiment, and for $m\in\{5,10\}$, we show the trade-off between the average objective value and empirical probability of constraint violation. The trade-offs for the LRO sets are obtained by applying the training algorithm for the $\eta$ values provided above, fine-tuning the size of the uncertainty set, and reporting the best objective values for a range of $\hat{\eta}$ values. We observe lower objective values for the LRO sets across the relevant levels of constraint violation. Note that in general, at the same out-of-sample objective value, the mean-variance RO and nominal Wasserstein DRO sets suffer from a much higher empirical probability of constraint violation. Therefore, the decision-focused uncertainty sets are more effective than uncertainty sets based on the distribution alone.
Comparing the LRO-RO and LRO-MRO sets, we note that the RO-based set has a higher variance in performance. However, as seen for $n=10$, it could outperform the MRO-based set. In terms of computation time, while the RO problems are less expensive to solve, the MRO framework (Wang et al.,, 2022) already presents a significant computational time reduction compared to Wasserstein DRO, especially when $K\ll N$.
In Figure 4, for $n=10$ stocks, a single $y$, and a single run of the experiment, we show the distribution of the optimal portfolio weights for all four uncertainty sets. The mean of the uncertain returns is $\mu=(0.5,0.55,0.57,0.59,0.65,0.67,0.7,0.71,0.77,0.86)$, sorted in ascending order (bottom to top in Figure 4), and the corresponding diagonals of $\Sigma$ are $(0.64,0.08,0.25,0.03,0.06,0.25,0.11,0.36,0.13,0.02)$. The covariances between the entries are omitted here for brevity. To mitigate the risks, when $\hat{\eta}$ is low the distribution of portfolio weights is either skewed towards the relatively lower variance stocks, as is the case for the the RO sets, or is close to an even distribution, as is the case for the DRO sets. Intuitively, one would invest more in stocks with higher returns and higher risks when the proxy of risk tolerance, $\hat{\eta}$, is high. Indeed, we observe that the relatively high-risk, high-return stock, the third-to-last stock (depicted in orange), receives larger investments as $\hat{\eta}$ increases. Note, however, that the reshaped sets, LRO-RO and LRO-MRO, are able to identify this profitable stock and increase its portfolio weight even at relatively low values of $\hat{\eta}$. By identifying a more compact region of uncertainty targeted for the problem family, they are able reduce conservatism in the portfolio selection process and take more profitable positions.
Figure 3: Portfolio. Average trade-offs between out-of-sample objective values and probability of constraint violation across 20 experiments. The shaded region represents the 0.3 to 0.7 quantiles. The LRO results are obtained through training procedures with $\eta$ values in the range $[0,0.3]$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Portfolio. Optimal portfolio weights chosen by the four methods for various $\hat{\eta}$. The colors represent the true mean $\mu$ of the returns.

### v5

@@ 4 Probabilistic guarantees of constraint satisfaction
In the previous section, we have proved convergence to a feasible solution of the sampled problem (13), which does not exactly correspond to the probabilistic guarantees of constraint satisfaction (11). To arrive at our desired guarantees, where the inner probability is with respect to the unknown joint distribution ${\mathbf{P}}_{(u,x)}$, we need to analyze the generalization behavior of the robust solutions on out-of-sample data. In addition to Assumption 3.4, we need a boundedness assumption on the function from the $\mathop{\bf CVaR}$ definition.
@@
The function $g(z(\theta,x),u,x)-\alpha)_{+}/\eta+\alpha$ maps to values in $(-c^{\rm min},c^{\rm max})$, where $0\leq c^{\rm min},c^{\rm max}<\infty$, for all $\theta,\alpha,u,x$ in its support.
Under Assumption 4.1, we then let $C=c^{\rm min}+c^{\rm max}$ and define the normalized function
$\psi_{w}(u,x)=\frac{1}{C}\left(c^{\rm min}+\frac{(g(z(\theta,x),u,x)-\alpha)_{+}}{\eta}+\alpha\right),$ (16)
where $w=(\theta,\alpha)\in\mathcal{Z}\subset{\mathbf{R}}^{n+1}$. Note that when Assumption 3.4 holds, we have $\mathcal{W}=\mathcal{D}_{\theta}\times\mathcal{D}_{\alpha}$ being a compact set. The function $\psi_{w}$ then maps to values in $(0,1)$, and we consider the entire function class
$\Psi=\{\psi_{w}\}_{w\in\mathcal{W}},$
@@
$\sup_{Q}N^{\text{cov}}(\delta,\Psi,L_{2}(Q))\leq\left(\frac{K}{\delta}\right)^{V},$ (17)
for every $0<\delta<K$. Then, when Algorithm 1 converges to an $\epsilon$-stationary point of (13), $\theta^{\star}$, we have the finite sample probabilistic guarantee
${\mathbf{P}}^{N}\left({\mathbf{P}}_{(u,x)}(g(z(\theta^{\star},x),u,x)\leq 0)\geq 1-\eta\right)\geq 1-\beta,$ (18)
where $\beta=(D\sqrt{N}C\tau/\sqrt{V})^{V}\exp(-2NC^{2}\tau^{2})$, $D$ is a constant that depends only on $K$, and $\tau\leq-(c^{\min}+\kappa+\epsilon)/C$.
@@
${\mathbf{P}}^{N}\left(\sup_{\psi_{w}\in\Psi}\left\|\frac{1}{N}\sum_{i=1}^{N}\psi_{w}(u^{i},x^{i})-{\mbox{\bf E}}_{(u,x)}[\psi_{w}(u,x)]\right\|\leq\tau\right)\geq 1-\left(\frac{D\sqrt{N}C\tau}{\sqrt{V}}\right)^{V}\exp(-2NC^{2}\tau^{2}).$
By the convergence of Algorithm 1 to an $\epsilon$-stationary point of (13), we have that the empirical $\mathop{\bf CVaR}$ will be close to $\kappa$. Given equation (16), this means that at $w^{\star}=(\theta^{\star},\alpha^{k})$, we have
$\frac{1}{N}\sum_{i=1}^{N}\psi_{{w}^{\star}}(u^{i},x^{i})\leq(1/C)(c^{\min}+\kappa+\epsilon).$
Therefore, as $\psi_{{w}^{\star}}\in\Psi$, we have
${\mathbf{P}}^{N}\left({\mbox{\bf E}}_{(u,x)}[\psi_{w^{\star}}(u,x)]\leq(1/C)(c^{\min}+\kappa+\epsilon)+\tau\right)\geq 1-\beta,$
@@
${\mathbf{P}}^{N}\left(\mathop{\bf CVaR}(g(z(\theta^{\star},x),u,x),\eta)\leq 0\right)\geq 1-\beta,$
which implies (18). ∎
This probabilistic guarantee holds for our data-driven solutions $z(\theta^{\star},x)$. We note that the covering number requirement holds when $\Psi$ is a class of Lipschitz functions, which is implied when the constraint function $g$ is Lipschitz in $z$, and $z$ is Lipschitz in $\theta$. For instance, when $g$ is affine or a maximum of affine functions in variable $z$, it is Lipschitz in $z$ for all $(u,x)\in\mathcal{D}_{(u,x)}$,  i.e.,
$|g(z,u,x)-g(z^{\prime},u,x)|\leq\bar{L}\|z-z^{\prime}\|_{2}\quad\forall z,z^{\prime}\in\mathcal{Z},~{}(u,x)\in\mathcal{D}_{(u,x)},$
with some Lipschitz constant $\bar{L}$. Then, if $z$ is Lipschitz in $\theta$, i.e.,
$|z(\theta,x)-z(\theta^{\prime},x)|\leq L_{z}\|\theta-\theta^{\prime}\|_{2}\quad\forall\theta,\theta^{\prime}\in\mathcal{D}_{\theta},~{}x\in\mathcal{D}_{x},$
we have that $g$ is Lipschitz in $\theta$ with Lipschitz constant $\bar{L}L_{z}$. Then, $\psi_{w}\in\Psi$ is Lipschitz in $w=(\theta,\alpha)$ with Lipschitz constant $L\leq(1/C)((\bar{L}L_{z}+1)/\eta+1)$. We can now derive the following covering number condition.
@@
When $\psi_{w}$ is $L$-Lipschitz in $w$ for all $u$ and $x$, we can apply standard bounds on the covering number of Lipschitz losses (van der Vaart and Wellner,, 1996; Pollard,, 1890; Namkoong and Duchi,, 2017). In this case, the covering number condition (17) reduces to
$\sup_{Q}N^{\text{cov}}(\delta,\Psi,L_{2}(Q))\leq N(\delta/L,\mathcal{Z},\|\cdot\|_{2})\leq\left(1+\frac{(\theta^{\rm max}+\alpha^{\rm max})L}{\delta}\right)^{q+1}\leq\left(\frac{K}{\delta}\right)^{q+1},$
for every $0<\delta<K$, where $K=\delta+(\theta^{\rm max}+\alpha^{\rm max})L$. Then, $\beta=(D\sqrt{N}C\tau/\sqrt{q+1})^{q+1}\exp(-2NC^{2}\tau^{2})$.
5 Hyperparameter tuning
[…]
Table 2: Default hyperparameter values and suggested ranges for the training algorithm 1.
$\rho$ $\alpha^{0}$ $\lambda^{0}$ $\mu^{0}$ $\gamma$ $\kappa$ $\sigma$ $\tau$ $k_{\text{max}}$ $t_{\text{max}}$ $\eta$ $|B|$
$[0.00001,5]$ 0 $[0.1,2]$ $[0.1,2]$ $[0,0.5]$ $-0.01$ 1.005 0.95 15 20 $(0,0.3]$ 200
Among the solutions obtained from the hyperparameter combinations, we use a validation dataset to select the one with the best performance. We prioritize grid-searching over the most impactful hyperparameters, such as the radius, training $\eta$ value, and the objective/constraint multipliers. For the other parameters, while the given default values may not be optimal, they work consistently well for our numerical experiments.
@@ 6 Examples
We run the following experiments on the Princeton Institute for Computational Science and Engineering (PICSciE) facilitywith 20 parallel 2.4 GHz Skylake cores. We solve the inner optimization problems (1) with the Clarabel optimizer with default settings. The learning procedure relies on the Python package cvxpylayers (Agrawal et al., 2019a, ), and the code to replicate the experiments (along with the baselines) can be accessed at
@@
We compare our method with the following baselines. To accommodate non-contextual methods, we restrict ourselves to discrete contexts, so the problems can be solved separately in each context. While we can also use clustering to restrict the number of representative contexts (for problems where the context does not appear in the formulation), here we use the exact contexts for a better comparison. For all experiments, we split the data into three components: training, validation, and testing. For all methods, we tune hyperparameters based on the performance on the validation dataset, for a validation probability of constraint violation below 10%, and for the best (lowest) 10% worst-case validation value among all sets that satisfies the criteria. We report the corresponding test-set values.
  • Learning for decision-focused RO (LRO) (our method). We apply the training method given in Section 3.2, for an ellipsoidal uncertainty set with $p=2$, and hyperparameter values from Table 2.
  • Mean-Variance RO (MV). We perform mean-variance RO, where the shape of the uncertainty set is determined by the empirical mean and variance of the data (Ben-Tal and Nemirovski,, 2000). We use grid search over the set size $\rho$ as hyperparameter tuning.
  • Contextual Mean-Variance RO (CMV). We perform contextual mean-variance RO, as described in Section 2.2, and use grid search over the set size $\rho$.
  • Wasserstein DRO (W). We perform Wasserstein-1 DRO (Esfahani and Kuhn,, 2018), where the uncertainty set is constructed around individual data points, regardless of the context. We again perform grid search over the set size $\rho$.
  • Wasserstein DRO with separate contexts (CW). We perform Wasserstein-1 DRO for each context, where the uncertainty set is constructed around individual data points for the particular context. We average the final result across all contexts.
  • Scenario Approach with separate contexts (CSA). We implement the scenario approach, inspired by Calafiore and Campi, (2005). For each context, we require the constraint to be satisfied by the associated training and validation uncertainty data. We average the results across all contexts.
  • End-to-End Conditional RO (ECRO). We implement the end-to-end approach given by Chenreddy and Delage, (2024), where a contextual neural-network is trained. We use grid search over $\eta\in[0.05,0.3]$ and $\gamma\in[0.1,0.9]$, which controls their objective $\mathop{\bf CVaR}$ and a conditional coverage objective, with 6 values in each interval. For other hyperparameters, we use the values given in their work.
  • Linear-Convex Ordering RO with separate contexts (CLCX). We implement the LCX uncertainty set given by (Bertsimas et al.,, 2018, Section 7), which is a custom constraint-based uncertainty set built using hypothesis testing. We use the hyperparameters $\alpha=\epsilon=0.1$, as given in their numerical experiments. We repeat the procedure for each context, and average the results across all contexts.
For each setting, we repeat the experiment 10 times, each with a new independent dataset, and evaluate the performance metrics in Table 3. For all methods with size tuning, we use $60$ values of $\rho$, and for our method, additionally grid over 6 values of $\eta$ and 6 combinations of $(\lambda^{0},\mu^{0},\gamma$) in the given ranges.
Table 3: Performance metrics, averaged over 10 runs.
Name Description $\hat{\eta}$ Test-set empirical probability of constraint violation Worst-case Test-set 10% worst-case objective value Average-case Test-set average-case objective value $\mathop{\bf CVaR}$ Test-set 10% Conditional Value-at-Risk
@@ 6.1 Portfolio optimization
We consider the classic portfolio management problem, where we select a portfolio $z\in{\mathbf{R}}^{n}$ of stocks to maximize $u^{T}z$, where $u\in{\mathbf{R}}^{n}$ are the uncertain asset returns. We assume to have access to contextual data on the market factors $F\in{\mathbf{R}}^{(n,2)}$, which influences the returns covariance, as well as data $h\in{\mathbf{R}}^{n}$ that influences the mean of the returns. This translates to the RO problem
$\begin{array}[]{ll}\mbox{minimize}&t\\ \mbox{subject to}&-u^{T}z\leq t\quad\forall u\in{\mathcal{U}}(\theta,x)\\ &\mathbf{1}^{T}z=1,\quad z\geq 0,\end{array}$
where the contextual information $x=(h,F)$ does not appear in the problem objective or constraints.
@@
We consider the cases where $n=10$ and $n=30$ stocks. For this problem, the uncertainty size $m=n$. We generate 20 different contexts, where for each context we have $x^{i}=(h^{i},F^{i}),$ with $h^{i}$ sampled uniformly on $[0.5,1]$ and the components of $F^{i}$ sampled independently from $N(0,1)$. The contextual uncertain returns $u^{i}$ is sampled from a multivariate normal distribution $N(\mu^{i}(h^{i})),\Sigma^{i}(F^{i}))$, where $\mu^{i}(h^{i})=0.7\mu+0.3h^{i}$, and $\Sigma^{i}(F^{i})=0.2F^{i}F^{iT}$. The nominal mean $\mu$ is sampled uniformly on $[0.5,1]$. We consider the cases with $N=2000$ and $N=1000$ samples, with the same number of samples for each context. For both cases, we use $30\%$ of the data for training, $20\%$ for validation, and $50\%$ for testing.
@@ Results.
In Table 4, we compare the performance of all methods, and note that our method outperforms the others in all objective metrics. This means that not only do we do better in the 10% worst-case and average case, the total loss we incur in the 10% worst cases is also consistently lower. The second-best method is CMV, where we already identify a good mapping from the context to the uncertainty set. Our learning method initializes from this set, and moves to less conservative mappings. As shown, these improvements cannot be found with traditional size-tuning. The other contextual method, ECRO, also performs relatively well, but suffers from the conservative coverage requirement, which reduces $\hat{\eta}$ in exchange for worse out-of-sample performance. The CSA uncertainty set performs the worst in terms of probability of constraint violation, as it generally requires more data than the RO methods that can tune a size parameter. Since the training data is consistent across all methods, it is unable to achieve the target guarantee.
The non-contextual MV and W uncertainty sets, which use all data-points regardless of the context, suffer from this generalization, as expected. However, such formulations are much more tractable than those of CSA, CLCX, and CW, for which a separate problem must be solved for each context. These methods also did not perform as well as the contextual methods, perhaps due to the (relatively) limited number of data-points available for each context. Furthermore, these methods could only be employed because we restricted the number of contexts; if we had unique contexts, the contexts may also need to be treated as an uncertainty set, or clustered, and might introduce additional conservatism in compensation for the information loss.
We note that the above behavior trends are consistent across the 4 scenarios, and comparing scenarios, a larger uncertainty set size $m=n$ (which for this problem indicates a larger selection of stocks) allows for better overall solutions, as does a larger $N$. The effects of these values are relatively similar for all methods, besides CSA, which is heavily impacted by the lack of data; it requires more data to achieve the 10% guarantee. We note that a larger $n$ may increase the solving and training time for all problems, as it increases the complexity. A larger $N$, on the other hand, might not affect the training time, as long as the batch size is kept consistent.
Table 4: Portfolio optimization comparisons averaged across 10 runs. The best value for each category is in bold.
$n=10$ $N=2000$ $N=1000$
Method $\hat{\eta}$ Worst-case Average-case $\mathop{\bf CVaR}$ $\hat{\eta}$ Worst-case Average-case $\mathop{\bf CVaR}$
LRO $0.0080$ $\bf{-0.818}$ $\bf{-0.862}$ $\bf{-0.751}$ $0.0094$ $\bf{-0.800}$ $\bf{-0.855}$ $\bf{-0.739}$
MV $0.0479$ $-0.578$ $-0.812$ $-0.441$ $0.0407$ $-0.573$ $-0.809$ $-0.440$
CMV $0.0469$ $-0.780$ $-0.823$ $-0.741$ $0.0434$ $-0.757$ $-0.809$ $-0.688$
W $0.0452$ $-0.551$ $-0.810$ $-0.403$ $0.0120$ $-0.553$ $-0.807$ $-0.415$
CW $0.0283$ $-0.621$ $-0.828$ $-0.550$ $0.0034$ $-0.479$ $-0.810$ $-0.403$
SA $0.1494$ $-0.489$ $-0.803$ $-0.338$ $0.2572$ $-0.484$ $-0.812$ $-0.344$
ECRO $0.0014$ $-0.656$ $-0.816$ $-0.530$ $\bf{0.0006}$ $-0.639$ $-0.822$ $-0.508$
LCX $\bf{0.0013}$ $-0.476$ $-0.810$ $-0.320$ $0.0056$ $-0.425$ $-0.820$ $-0.291$
$n=30$ $N=2000$ $N=1000$
Method $\hat{\eta}$ Worst-case Average-case $\mathop{\bf CVaR}$ $\hat{\eta}$ Worst-case Average-case $\mathop{\bf CVaR}$
LRO $0.0018$ $\bf{-0.902}$ $\bf{-0.925}$ $\bf{-0.880}$ $0.0082$ $\bf{-0.901}$ $\bf{-0.927}$ $\bf{-0.841}$
MV $0.0844$ $-0.721$ $-0.831$ $-0.656$ $0.0858$ $-0.723$ $-0.828$ $-0.659$
CMV $0.0410$ $-0.871$ $-0.898$ $-0.861$ $0.0320$ $-0.864$ $-0.894$ $-0.840$
W $0.0899$ $-0.668$ $-0.842$ $-0.558$ $0.0638$ $-0.666$ $-0.836$ $-0.555$
CW $0.0283$ $-0.621$ $-0.828$ $-0.550$ $0.0090$ $-0.610$ $-0.822$ $-0.562$
SA $0.2850$ $-0.652$ $-0.830$ $-0.549$ $0.4310$ $-0.598$ $-0.849$ $-0.488$
ECRO $\bf{0}$ $-0.773$ $-0.857$ $-0.697$ $\bf{0}$ $-0.758$ $-0.854$ $-0.673$
LCX $0.0001$ $-0.628$ $-0.820$ $-0.528$ $0.0007$ $-0.555$ $-0.834$ $-0.447$